### v2

Abstract
This article presents the automatic checking of research outputs package ACRO, which assists researchers and data governance teams by automatically applying best-practice principles-based statistical disclosure control (SDC) techniques on-the-fly as researchers conduct their analyses. ACRO distinguishes between: research output that is safe to publish; output that requires further analysis; and output that cannot be published because it creates substantial risk of disclosing private data. This is achieved through the use of a lightweight Python wrapper that sits over well-known analysis tools that produce outputs such as tables, plots, and statistical models. This adds functionality to (i) identify potentially disclosive outputs against a range of commonly used disclosure tests; (ii) apply disclosure mitigation strategies where required; (iii) report reasons for applying SDC; and (iv) produce simple summary documents trusted research environment staff can use to streamline their workflow. The major analytical programming languages used by researchers are supported: Python, R, and Stata. The ACRO code and documentation are available under an MIT license at https://github.com/AI-SDC/ACRO
{IEEEkeywords}
Data privacy, data protection, data science, privacy, Python, R, Stata, statistical disclosure control.
1 INTRODUCTION
Statistical agencies and other custodians of secure facilities such as trusted research environments (TREs) [1] provide researchers with access to confidential data under the ‘Five Safes’ data governance framework [2]. This enforces five orthogonal layers of safety procedures with the last requiring explicit checking of research outputs for disclosure risk. This can be a time-consuming and costly task, requiring skilled staff. This article discusses the development of a free and open source tool for automating the statistical disclosure control (SDC) of routine research outputs such as tables, plots, and statistical models. The goal is to make the clearance process more efficient and timely, and to allow the skilled checkers to focus their attention on the less straightforward cases.
More specifically, ACRO (automatic checking of research outputs) assists researchers and output checkers by distinguishing between research output that is safe to publish, output that requires further analysis, and output that cannot be published because of substantial disclosure risk. This is achieved through the use of Python’s capacity to override standard commands for creating tables, regressions, and other queries. This keeps the syntax identical while simultaneously augmenting analysis commands by running and reporting appropriate disclosure risk assessments. A schematic illustration of the ACRO workflow is shown in Figure LABEL:fig:schematic.
The operational design philosophy is extensively documented by Green et al. [3] who studied the characteristics that an automated SDC solution needs to have to be feasible, effective, and a positive choice for users. The essential criteria are that it should be:
  1. 1.
    Free and open source: ACRO is available on GitHub under an MIT License at: https://github.com/AI-SDC/ACRO.
  2. 2.
    Easy to obtain: in addition to the source code available on GitHub, a Python package is available through the Python Package Index (PyPI), which provides a simple management system for distributing Python libraries and facilitating continuous updates. Similarly, the R ACRO package [4] is available through the Comprehensive R Archive Network (CRAN). ACRO is cross-platform, supporting GNU/Linux, macOS, and Windows; with v0.4.6 supporting Python 3.8–3.12.
  3. 3.
    Easy to use: ACRO reduces barriers to adoption via a front-end application programming interface (API) that is identical to those already commonly used by researchers in their favoured language. Researchers prepare their data and statistical queries in the usual way, in their preferred language. When they are satisfied with the format of a query that generates, for example, a table or a regression, they then repeat it using the same commands prefixed by acro. The lightweight translation functions then call the Python back-end, which executes the queries and performs the requisite output checks. Easy to understand help and documentation is available within the GitHub repository.
  4. 4.
    Comprehensive: ACRO identifies potentially disclosive outputs with a range of commonly used disclosure tests, applies mitigation strategies where required, and flags outputs for human review that cannot be definitively classified as safe or unsafe. The current version of ACRO uses suppression and rounding to protect privacy, which are the most common forms of disclosure mitigation strategy.
  5. 5.
    Transparent: applying disclosure mitigation where required, but with clear and rapid explanations that enable (i) researchers to understand and revise their requests; and (ii) an auditable record for TRE staff with simple summary reports to streamline their workflow; ACRO provides researchers with immediate feedback by displaying the results of both the checks and the output to the researcher, with the option for automatic suppression where appropriate. Details of the queries and results are stored in a list, which may subsequently be written to file for review by a human output checker. ACRO gives researchers control over the outputs that are submitted for review, e.g., the removal of unwanted outputs and choice of output format; currently JSON [5] or Microsoft Excel®.
  6. 6.
    Able to support exceptions under principles-based regimes; ACRO enables researchers to add comments to requested outputs and ensures that researchers supply supporting evidence for all outputs that cannot be classified as safe.
  7. 7.
    Consistent: providing the same results across different studies within a TRE, and across TREs; ACRO contains a single back-end code base constituting a single source of truth for performing checks with extensibility for different languages and ongoing support and consistency.
  8. 8.
    Scalable over users and outputs. Since ACRO is lightweight, session-based, and runs on the machine that generates outputs, it scales easily to many users and outputs.
  9. 9.
    Able to implement an organisation’s business rules for primary and secondary disclosure, which may vary across datasets or users. Within ACRO, SDC rules and parameters can be uniquely configured through a simple YAML [6] file specifying the risk appetite.
    The current version of ACRO does not yet address secondary disclosure (such as checking for differencing across tables), for two reasons. First, business rules for secondary checking are often not clear or comprehensive. Second, ACRO works by intercepting commands and assessing disclosure risk at the time the output is being produced. Analysing results post-hoc is a considerably harder problem, requiring the researcher to produce a lot more information and also locate the other outputs to be compared.
ACRO is a product of the “Semi-Automated Checking of Research Outputs” (SACRO) project funded by Data and Analytics Research Environments UK (DARE UK). [^1] https://dareuk.org.uk/driver-project-sacro/ This work builds upon a previous Eurostat-funded project, in which Green et al. [7] developed a proof-of-concept prototype for the proprietary Stata software. As part of the SACRO project, the SACRO team committed to review and re-develop the theory and operational guidelines for output SDC. The aim was threefold; first, to bring together key points from the SDC literature (and fill in some of the theoretical gaps) to provide an integrated guide to both theory and practice of output checking; second, to develop a new approach to output SDC based on classifications into groups; for details, see Derrick et al. [8]; third, to explicitly link theory to operational rules and their implementation in manual and automatic checking regimes.
Five UK TREs (OpenSAFELY [^2] https://www.opensafely.org at the University of Oxford and four Scottish Safe Havens) were involved in the SACRO project to provide detailed feedback on user and output checker perspectives; OpenSAFELY also took the lead in the design of a graphical user interface (GUI). This group also directly tested the feasibility of installing and allowing the Python code to run on their systems as TREs differ in their perceptions of Python’s ‘riskiness’. Moreover, the SACRO team contacted a large number of TREs in the UK and abroad, and set up a network of interested parties willing to be testers. Several engagement events with this group identified how they worked and what they would expect from an automatic solution.
Subsequently, a ‘Community of Interest’ group has been established to provide on-going peer support and new feature prioritisation. Currently, the mailing list333mailto:sdc-reboot@jiscmail.ac.uk has 50 subscribers from 25 different organisations. To illustrate the breadth of uptake, just some of those who are deploying SACRO and encouraging its use by researchers include: eDRIS/Public Health Scotland and the Scottish National Safe Haven; Clinical Practice Research DataLink (Medicines & Healthcare Products Regulatory Agency, England), and the Secure Anonymised Information Linkage (SAIL) Databank (Wales).
The remainder of this article is organised as follows. Section 2 describes the current state-of-the-art for output SDC with the Five Safes framework and SDC implementations. Section 3 details the SDC checks implemented with ACRO, as well as how to install and use the Python, R, and Stata packages. Section 4 describes the GUI for viewing and managing the results of SDC produced by ACRO. Finally, Section 5 discusses future plans.
2 BACKGROUND
2.1 FIVE SAFES AND OUTPUT DISCLOSURE CONTROL
The Five Safes framework [9] is a set of principles that enable services to provide safe research access to their data and has been adopted by a range of TREs, including the UK Office for National Statistics (ONS), Health Data Research UK (HDR UK), and the National Institute for Health Research Design Service (NIHR), as well as many others worldwide.
The Five Safes are composed of:
  • Safe data: (pseudo)anonymisation and minimisation of data provided to researchers.
  • Safe projects: research projects have been approved and make appropriate use of data.
  • Safe people: researchers trained and authorised to use data safely.
  • Safe settings: a secure environment that prevents unauthorised use and ingress or egress of data other than through approved channels.
  • Safe outputs: screened and approved outputs that are non-disclosive.
TRE staff commonly follow the SDC handbook by Griffiths et al. [10], which provides advice on how to assess specific statistical outputs within the context of the Five Safes, as well as advice to organisations about how to set up SDC systems. Additionally, Jefferson et al. [11] provide recommendations for the SDC of machine learning models trained within TREs. However, ensuring ‘safe outputs’ is a complex and often costly human labour-intensive process. As DARE UK [12] have recently highlighted: “there are already significant issues with staffing resources to support statistical disclosure control (safe outputs). This staffing issue is acting as a barrier to scaling up the use of TREs.”
2.2 SAFE VS. UNSAFE STATISTICS
Ritchie [13] first introduced the concept of ‘safe statistics’, which aims to classify statistics as ‘safe’ or ‘unsafe’. ‘Unsafe’ statistics, such as tables of frequencies, present a number of potential disclosure risks. ‘Safe’ statistics in contrast, such as the coefficients of correlation from model estimates, have no effective risk.
Mitigation strategies commonly suppress, round, or add noise to ‘unsafe’ outputs in order to prevent disclosure [14]. For example, a minimum threshold rule applied to the number of observations used by a statistic ensures that there is sufficient uncertainty with respect to any individual respondent. Dominance rules protect large respondent values from being approximated where the contribution to a statistic is dominated by only a few individuals. For example, the $p$%-rule sorts the $N$ observations by magnitude and checks whether the sum of the smallest $N-3$ observations is at least $p$of the largest observation. The $NK$ rule checks that the largest $N$ observations contribute less than $K$of the total. Not all aggregation statistics are considered safe; typically, reporting minima or maxima values of a subgroup are prohibited.
Regressions are generally protected by checking that the residual degrees-of-freedom exceeds a minimum threshold. The purpose is to prevent an equation from masquerading as a model. Residual degrees of freedom is defined as the number of observations minus the number of restrictions embodied in the test or model. For example, for a simple linear regression it is $N-K$, where $K$ is the number of coefficients including the intercept; for chi-square it is $N$ test degrees of freedom. Following Brandt et al. [15], 10 is typically used as the minimum.
Devising rules for each individual statistic used by analysts is infeasible, however it is possible to combine statistics into groups based not on statistical relation, but on common disclosure risks and solutions. Consequently, Ritchie et al. [16] propose the use of ‘statistical barns’ (statbarns) as a means to classify statistics for disclosure control purposes. A statbarn is a collection of statistics that share the same characteristics for disclosure control purposes. That is, their mathematical form is similar; they share the same risks; they share the same responses to those risks; and output checking rules are applicable to all. For example, pie charts, histograms, and scatter plots are all forms of frequency table. Table LABEL:table:statbarns lists the 14 statbarns that have been identified, with 12 classified for output checkers.
2.3 PRINCIPLES-BASED DISCLOSURE CONTROL
Traditional approaches to SDC within TREs have focused on rules-based approaches, where a rule is a hard limit with no exceptions permitted. More recently, it has been recognised that a principles-based approach to SDC is more effective since defining hard rules that are neither too restrictive nor too loose is extremely difficult [17]. For example, a frequency table with zeros in some cells is highly disclosive if they reveal that no one in a given region earned more (or less) than $k$ USD. In contrast, reporting that no one born without organ $k$ was diagnosed with cancer of said organ would usually be recognised as a “structural zero”, and permitted.
In principles-based SDC, both researchers and output checkers undertake SDC training. They are provided with a set of heuristics rather than strict, definitive rules. These heuristics serve as initial guidelines. If a researcher wishes to request approval for outputs that deviate from these guidelines, they must meet three conditions: (i) ensure the outputs do not disclose sensitive information; (ii) demonstrate the outputs’ significance; and (iii) justify why the request is exceptional. The responsibility lies with the researcher to demonstrate that any potentially risky outputs do not disclose sensitive information; although the final decision rests with the checker. Since there are no strict rules, both the researcher and the checker need expertise in understanding disclosure risks and exercising judgement.
Semi-automated output checking therefore aims to improve the rigour and consistency of the output disclosure control process and reduce human workload by automatically applying best-practice principles-based SDC techniques. That is, automatically identifying, reporting, and optionally applying mitigation strategies to disclosive outputs where possible. Outputs are categorised as ‘safe’ or ‘unsafe’. ‘Safe’ outputs requiring no or minimal further changes can be expedited through the clearing process, whereas ‘unsafe’ outputs can be prioritised for human review [13].
The use of semi-automated SDC tools for checking research outputs therefore enables:
  • An explicit change in the dynamics so that SDC is something done with researchers rather than to them.
  • Reduced costs for TREs.
  • Improved efficiency with skilled personnel focusing on the areas of most significant risk.
  • Improved security and public confidence by reducing the chance of human error.
  • Improved auditing by generating outputs in a structured format.
  • Improved user experience by making output checking quicker.
  • Increased consistency of decision making within and between TREs.
2.4 TOOLS FOR SDC
A small number of SDC tools have been produced to assist in the process of achieving ‘safe outputs’, such as sdcTable [18] and tauArgus [19] by Statistics Netherlands; however they require expert knowledge of SDC to use effectively. Moreover, they are exclusively designed for tabular (frequency and magnitude) outputs, and do not cover the range of statistics produced by researchers. They are predominantly used for regularly repeating reports in National Statistics Institutes. Metadata must be created to describe each output and control parameters configured. The need to rewrite the metadata for each table makes these tools poorly suited for research use. However, for scenarios where the same tables are repeatedly generated and secondary differencing is considered a significant problem, the investment in setting up the tools can be cost-effective.
DataSHIELD [^4] https://datashield.org is an infrastructure and suite of R packages [20] that aims to enable remote and non-disclosive analysis of distributed sensitive research data while avoiding the normal practice of human involvement in output checking. It operates under the principle that researchers never see or have access to the underlying data, even for manipulation. This is achieved by providing a set of restricted bespoke commands for performing data manipulation and querying. This approach implements rules-based rather than principles-based SDC and some rules (e.g., allowing min/max values, and not suppressing table cells with zero counts) differ significantly from the standard practice as described in the SDC handbook [10].
The Research Data and Service Centre (RDSC) of the Deutsche Bundesbank provide two tools to allow researchers to check whether the results generated in their research project comply with the RDSC’s SDC rules [21]. Researchers run a special command directly after generating their results and get immediate feedback. The RDSC provide nobsdes5/nobsreg5, which are a set of Stata ado files, and the sdcLog [22] package, which provides a set of new R functions. These tools help researchers in checking descriptive statistics and models, and in calculating extreme values that are not individual data; however these packages have not been updated since March 2022. Since these tools implement new commands, researchers must learn them and adapt their workflow.
With the aim of developing a more general semi-automated solution for use by researchers in TREs, improving the efficiency of the process, and (where applicable) reducing the amount of user training required, a recent Eurostat project [7] developed a proof-of-concept prototype in Stata where primary disclosure is regulated by a set of simple rules. Here we build upon the lessons learned from developing that prototype with the aim of expanding its coverage and impact whilst retaining a minimal learning curve. We achieve this by adopting the principles-based approach to SDC outlined by Ritchie et al. [16] and develop the ACRO toolkit, which is primarily implemented in Python, but with cross-language support for R and Stata, using analytical interfaces familiar to researchers.
3 THE ACRO TOOLKIT
3.1 DISCLOSURE CONTROL IMPLEMENTED
For tabular data (e.g., cross tabulation and pivot tables), ACRO prohibits the reporting of the maximum or minimum value in any cell that represents a sub-group of one or more contributors. Moreover, ACRO suppresses, and reports the reason why, the value of an aggregation statistic (mean, median, variance, etc.) for any cell is deemed to be sensitive. The current version of ACRO supports the three most common tests for sensitivity: ensuring the number of contributors is above a frequency threshold, and testing for dominance via $p\%$ and $NK$ rules. In addition, any outputs found to include negative or missing values are automatically flagged for human review since the results of SDC are not well defined in these circumstances.
For tables, ACRO builds a series of suppression masks, which indicate which cells are to be suppressed for each check. A summary outcome table indicating which suppression rule was applied to each cell is presented to the researcher, along with the result of the query.
For regressions such as linear, probit, and logit, the tests verify that the number of residual degrees of freedom exceeds a threshold. There is currently limited support for graphical plots, however ACRO supports the SDC of histograms using the same rules as for frequencies since these fall under the same statbarn category; if a histogram is found to be disclosive and automatic suppression is enabled, the plot is hidden from the researcher. Additionally, survival plots such as Kaplan-Meier are protected using a rounding mitigation strategy.
As noted above, all of these tests and checks are configurable according to the TRE’s risk appetite. The data custodian, e.g., TRE staff member, specifies the parameter values used for the output checks in a YAML configuration file, which is loaded upon ACRO initialisation. The default ACRO parameters are shown in Table LABEL:table:params. Future releases will offer the option to override these on a dataset, or even attribute level.
3.2 PYTHON PACKAGE
Python is a popular multi-platform language widely used for data analysis and machine learning. PyPI provides a simple and easy to use package management system for distributing open source Python libraries. Pandas [23] and the statsmodels [24] Python libraries are mature, popular, and well-supported packages for data analysis, statistical testing, and statistical data exploration. Pandas is used by more than 55% of all Python users [25]. The use of Python as the primary implementation therefore enables the leveraging of existing expertise and community support with these packages so that the ACRO front-end can be as similar to the API researchers already know and trust, and further facilitates the rapid development of disclosure checking functionality on the back-end. As the PyPI distribution system is simple and allows the use of semantic versioning, it supports a rapid and iterative develop-and-deploy strategy to provide continuing functionality and improvements. The Python ACRO package also depends on Matplotlib [26] to support data visualisations, NumPy [27] as a core numerical matrices package, and PyYAML [28] for YAML configuration file support.
The newest version of ACRO may be installed from PyPI:
\$ pip install acro
Then, to get started, instantiate an ‘ACRO’ object as seen in Figure LABEL:fig:pyinit, which is used to store and maintain outputs. The ‘ACRO’ constructor optionally accepts a string parameter config that can be used to specify an alternative YAML rule configuration file, and a Boolean parameter suppress that specifies whether automatic suppression should be applied within the session.
Standard data analysis operators can then be used by pre-fixing the function with ‘acro’. Figure LABEL:fig:crosstab shows how the ‘pandas.crosstab’ function can be used with the same arguments and in the same way, operating on a ‘pandas.DataFrame’, df. However, in this instance, the R/G column and a single cell in column N is automatically suppressed since it violates disclosure checks. The output also includes an explanation of which disclosure checks failed for each cell, a summary of the checks, and the name assigned to the output so that it can be referenced subsequently; for example to add comments or request an exception from a human checker.
Similarly, pivot tables can be created as in Figure LABEL:fig:pivot. Note that the only API difference is calling ‘acro.pivot_table’ instead of ‘pandas.pivot_table’. In this example, the entire column R/G has been withheld from the researcher as a result of SDC. Since the tables in these examples failed SDC, exception requests must be provided, such as as in Figure LABEL:fig:exception.
In addition to common Pandas tabular operations, ACRO provides disclosure control functions for common statsmodels regression operations. Figure LABEL:fig:ols shows an example where an ordinary least squares is performed on $x$ and $y$ where the residual degrees of freedom are checked. Note that the only API difference is calling ‘acro.ols’ instead of ‘statsmodels.api.OLS’.
Limited support for graphical plots is available. For example, Figure LABEL:fig:hist shows the result of attempting to produce a disclosive histogram. The plot is therefore hidden from the researcher. Note that the only API difference is calling ‘acro.hist’ instead of ‘pandas.hist’.
The ‘ACRO’ class includes a number of session management functions to display, rename, remove, or add comments to existing outputs, so that the researcher maintains control over the outputs they wish to request. An example renaming an output for convenience is shown in Figure LABEL:fig:rename; an example adding comments to existing outputs is shown in Figure LABEL:fig:comments; and an example removing an unwanted output can be seen in Figure LABEL:fig:remove.
Moreover, as previously shown the researcher can request an exception for a particular output and include their supporting evidence; unsupported outputs can be added with the add_custom_output function and by specifying the path to any file. These unsupported outputs are then included in the final requested release and highlighted for the attention of a human checker. When the researcher is finished with their analysis, they simply call the finalise function:
$>$ acro.finalise()
If any of the outputs have failed the SDC checks at this stage and an exception has not already been requested, the researcher is prompted to provide one. An SDC report is then created in JSON and all outputs are written to a directory named “outputs” by default; a subdirectory with SHA-256 checksums [29] is also generated. Optionally, an alternative directory may be requested by supplying a string parameter path to finalise. Additionally, the report can be written in Excel® format by passing a string parameter ext set to xlsx. Having finalise run automatically when the object is destroyed (i.e., session closes) was considered, but for now this was decided against since researchers may choose to save their work between sessions and not wish to go through the full release process.
The automatic suppression of outputs can be enabled or disabled via the Boolean suppress constructor parameter. Regardless of whether suppression is applied, the results of the checks are included in the final JSON report. This report file can then be ingested by the GUI described in Section 4 to display the disclosure results and suppression can be applied on a case by case basis.
Within the ACRO Python package, the functionality of the ‘ACRO’ class is split into a number of separate classes for maintainability and extensibility. A ‘Tables’ class contains the code necessary to perform disclosure checks on tabular data, such as crosstab. A separate ‘Regression’ class contains the code for checking regressions such as logit and probit. The ‘ACRO’ class thus inherits from each of these classes so that the researcher can invoke all of the functions directly on the instantiated object, which tracks the outputs generated and writes them upon finalise.
As Internet access is restricted from within TRE (virtual) environments, the ACRO package has extensive docstrings to support Python’s inbuilt help mechanisms; for example:
$>$ help(acro.ACRO)
Standard Python coding and naming practices have been used throughout [30, 31]. GitHub continuous integration (CI) runners automatically generate and publish API documentation using the Python docstrings written in numpydoc [32] format. The project also contains a configuration file for the pre-commit [33] tool, which is used to run mypy [34] type checking and enforce standards and code formatting with Ruff [35]. A GitHub bot is configured to automatically run pre-commit when any pull requests or pushes are made to the main GitHub branch. Additionally, a CI runner is configured to perform static code analysis with pylint [36].
Extensive pytest [37] unit tests have been written, which currently cover $>99$% of the Python source code. A CI runner is configured to automatically execute these tests and report the coverage results for all pull requests and pushes made to the main branch.
Moreover, CI runners are configured to pip install the main branch and execute the tests with a range of Python versions on Ubuntu, macOS, and Windows operating systems for quality assurance whenever a pull request has been merged to the main branch. An additional CI runner is used for building and publishing ACRO to PyPI; and we have recently added the ability to generate artifact attestations [38] to enable users to confirm the provenance of their ACRO packages.
Some Jupyter notebooks [39] demonstrating example code usage and output are available within the ACRO project repository. The currently implemented methods (v0.4.6) are listed as follows; for more details see the ACRO API documentation built with the tool Sphinx [40], which is available: https://ai-sdc.github.io/ACRO/.
3.2.1 PYTHON TABLE METHODS
3.2.2 PYTHON REGRESSION METHODS
3.2.3 PYTHON SESSION MANAGEMENT METHODS
3.3 R PACKAGE
The R ACRO package [4] is an example of cross-language support and is available through CRAN. It provides a set of R wrapper functions that execute Python back-end checking within a virtual environment via the reticulate package [41], which provides automatic conversions for types such as the R ‘data.frame’ to Pandas ‘DataFrame’. The source code is available under MIT license at: https://github.com/AI-SDC/ACRO-R. GitHub CI runners automatically validate the package for CRAN by running R CMD check [42] using different operating systems and Python versions to ensure cross-platform operability. CI runners also generate and publish API documentation using pkgdown [43] and generate test code coverage reports.
For regressions, the common R lm and glm functions were shadowed with equivalent versions implemented as acro_lm and acro_glm, respectively, exploiting the fact that statsmodels already supports R-style specification of formulae. For tabular data, the standard R table command has been shadowed with acro_table to perform cross tabulation analysis. While the dplyr package [44] is commonly used, no simple pivot table functions are provided; instead various combinations of groupby and summarize etc. are used. Therefore, at this stage of development, the Python cross tabulation and pivot table functions were directly interfaced with acro_crosstab and acro_pivot_table.
To install the R ACRO package from CRAN:
$>$ install.packages("acro")
Then to instantiate an ‘ACRO’ object:
$>$ library("acro")
$>$ acro_init(suppress = TRUE)
Subsequently, the Python package functions may be used, however for simplicity the package provides R helper functions that use underscore naming. For example, the crosstab function can be used identically to the Python version by calling acro_crosstab. Similarly, the pivot_table function becomes acro_pivot_table, etc. See the following list of R helper functions available; see the ACRO R reference manual on CRAN [^5] https://cran.r-project.org/web/packages/acro/acro.pdf for further details.
3.3.1 R TABLE METHODS
3.3.2 R REGRESSION METHODS
3.3.3 R SESSION MANAGEMENT METHODS
3.4 STATA INTERFACE
The ACRO Stata interface is implemented for versions below 15 and for versions 17 and above when the syntax changed. It makes extensive use of Stata’s SFIToolkit [45] to manage a Python session, transfer data in memory from Stata to a Pandas ‘DataFrame’ in the Python session, and results back to the Stata window.
A simple ‘acro.ado’ file defines a new function acro which takes as parameters either one of the ‘ACRO’ session management methods (adding init to start a session) or the name of a standard Stata function such as table, regress, etc. Stata’s inbuilt parsing functions are used to separate out the parts of command and pass them as lists to a Python function parse_and_run which handles the rest of the translation between the two languages.
The Stata ado files implementing the ACRO interface can be found in the ‘stata’ folder located within the main ACRO GitHub repository. [^6] https://github.com/AI-SDC/ACRO/tree/main/stata The files should be placed in the standard Stata ado file directory and the Python ACRO package must be installed, which includes a Stata to Python parser.
4 GRAPHICAL USER INTERFACE
OpenSAFELY have developed a platform-independent stand-alone tool for output checkers to view the outputs produced by ACRO, understand the risks associated with each output, make audited decisions, and produce zipped packages of files for release.
The viewer consists of (i) a Django [46] web application with a vanilla JavaScript user interface that renders a set of ACRO outputs for review; and (ii) an Electron [47] application and installer that bundles the web application. It is available on GitHub under GNU General Public License v3 at https://github.com/AI-SDC/SACRO-Viewer.
The application is packaged using standard tooling, e.g., as an msi for Windows, deb for Debian/Ubuntu, and dmg for macOS operating systems; all dependencies are bundled with the installers, including a chrome-based web browser [48], Python, and JavaScript runtimes. The Cypress [49] Javascript framework is used to implement a number of automated user interface tests for CI.
The viewer supports and renders a range of different file types for results from queries not yet supported by ACRO. Hence the viewer can be used for making and recording decisions, even if the researcher has not used ACRO during their analysis; automated disclosure risk analysis is not provided in those cases.
To view outputs, the user opens a directory containing the ACRO output files. The viewer then detects if there is ACRO generated metadata and uses that to display the files. If no ACRO data is available, it automatically generates it, adding each file in the directory as a custom output. The GUI viewer includes the ability to view relevant source code snippets and syntax highlighting to aid human review.
A screenshot of the GUI can be seen in Figure LABEL:fig:gui. The left pane displays the list of outputs to review. The top right pane provides an option to view the ACRO configuration specifying the risk profile. The centre right pane shows details of the currently selected output, including comments from the researcher, any exception requests, and the ACRO recommendation.
Moreover, the GUI provides buttons for the TRE checker to approve or reject the output and supply feedback, e.g., if the checker wishes to override the ACRO recommendation. The bottom right pane displays the currently selected output with any specific components that violate SDC checks highlighted in red. When checking has been completed, the top right ‘Release and download’ button can be used to record the overall comments and create a (zip) package for release, which automatically excludes any rejected outputs.
5 FUTURE PLANS
Future plans include ongoing development of both the main Python ACRO library and additional translation languages such as R and Stata to enhance functionality, support more analyses and respond to inevitable issues as they arise from an increased volume of testing. Future plans also include the production of an easy to install package for Stata; and the development of a toolkit that can be embedded within projects such as OpenSAFELY and medical TREs.
Future plans include implementing rounding mitigation strategies for frequency tables, improved support for graphical plots, and the ability to configure rules at a dataset or even attribute level.
Currently, the results of SDC checks for tables are presented by displaying the whole table, where the reasons for failure are presented within each cell; while this works well for small to medium size tables, it can be overwhelming for larger tables. Future versions may therefore present the row/column numbers of the offending cells with an explanation of the SDC check instead of the entire table.
Additional features and improved user experience will be facilitated by the further involvement of end-users and output checkers. Future versions may also extend coverage of more of the R Tidyverse [50] family such as dplyer, etc.; and rewriting parts of the translation to support different Stata versions.
As previously mentioned, the current version of ACRO does not implement secondary disclosure checks. Work is underway investigating how to at least flag potential differencing risks across the set of outputs from a research session, and in future, create a library of outputs which might allow secondary disclosure to be assessed, even if only partially.
Moreover, future versions of the GUI viewer may include outputs from checking trained machine learning models, such as those from the sacroml [51] Python package.
As the TRE infrastructure landscape evolves, and new lightweight dashboard frameworks such as Streamlit [52] increase adoption, the question of choice of tools for implementing the viewer design inevitably remain open.
The GUI viewer is currently intended for checkers to review the results produced by researchers, but not for researchers themselves to view, modify, and supply information for checkers; therefore, future plans include the development of such a GUI for researchers.
ACKNOWLEDGMENT
The authors are grateful to Elizabeth Green and Felix Ritchie for comments as this work progressed.
[…]
IEEEexample:BSTcontrol